OpenStudy (this):

what is the vertex of the function ? f(x)=(2x+2^2)+4 a,(2,4) b,(2,-4) c,(-2,4) d,(-2,-4)

OpenStudy (anonymous):

check your equation

OpenStudy (dumbcow):

to find vertex of quadratic equation: \[y=ax^{2} +bx +c\] use \[x = \frac{-b}{2a}\] then plug that into equation to get y

OpenStudy (this):

@NotSObright i fix it now thnks by the way

OpenStudy (dumbcow):

i think what you meant to write is \[2(x+2)^{2} +4\] ?? in which case its in vertex form already (x+2) means shift left 2 ... x=-2 +4 means shift up 4 ... y=4 vertex is (-2,4)

OpenStudy (this):

yes @bumbcow thanks